# Post-detection inference for sequential changepoint localization

### v2

@@ 2.1 A first confidence set
To test $H_{0,t}$, we draw $B$ many independent streams of data under $H_{0,t}$. For the $j$th sequence, denoted as $\{X_{n}^{j}\}_{n}$, we simulate data points until a change is detected at the stopping time $\tau_{j}^{\prime}$ using $\mathcal{A}$ or until time $L$, whichever occurs first. Then, we compute the same test statistics truncated at $L$ (which we discuss in detail in the next subsection), based on $\tau_{j}^{\prime},L$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$, which is denoted as $M^{j}_{t,L}$. We reject $H_{0,t}$ if $M_{t}>\operatorname{Quantile}(1-\alpha;M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$, and the confidence set is the collection of all $t\in\{1,\cdots,\tau\}$, for which we fail to reject $H_{0,t}$, i.e,
@@
$\mathcal{C}=\Big{\{}t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})\Big{\}}.$ (2.2)
Algorithm 1 provides an overview of the method.
Input: $\alpha\in(0,1)$, $B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$, algorithm $\mathcal{A}$ and data $\{X_{n}\}_{n}$ until a change is detected at $\tau$ using $\mathcal{A}$
Output: Confidence set $\mathcal{C}$ for the changepoint $T$
1 Detect change at $\tau$ using $\mathcal{A}$; Set $\mathcal{C}=\varnothing$
2 for $t=1,\cdots,\tau$ do
3       Compute some test statistic $M_{t}$ (e.g., as defined in (2.9)) based on $X_{1},\cdots,X_{\tau}$
4       for $j=1,\cdots,B$ do
5             Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+2}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{1}$, until a change is detected at $\tau_{j}^{\prime}$ using $\mathcal{A}$, or until time $L$, whichever happens first.
6             Subroutine $\mathcal{L}$: Compute $M_{t,L}^{j}$ (defined in (2.3),(2.4)) based on $\tau_{j}^{\prime}$, $L$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$
7
8       end for
9      if $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$ then
10            $\mathcal{C}=\mathcal{C}\cup\{t\}$
11       end if
12
13 end for
@@ 10            $\mathcal{C}=\mathcal{C}\cup\{t\}$
2.2 Subtleties regarding Subroutine $\mathcal{L}$
The problem with $L=\infty$ is that the stopping time may never be reached for the $j$-th sampled sequence, i.e., we may have $\tau^{\prime}_{j}=\infty$. One can set $L=\infty$ if $\mathcal{A}$ (i.e., its stopping time) satisfies the condition $\mathbb{P}_{T}(\tau<\infty)=1$, for all $T\in\mathbb{N}$ (if there is a change, $\mathcal{A}$ stops almost surely). This is a very mild assumption because any reasonable change detector will always eventually raise an alarm if there is a change. Indeed, all algorithms with finite ARL satisfy the condition, and indeed typically so do algorithms that control PFA. To avoid making even this mild assumption, one can simply pick $L<\infty$.
If $L<\infty$ and the $j$th sequence stopped because $\mathcal{A}$ raised an alarm at or before $L$, then we define $M^{j}_{t,L}$ as usual, but if $\mathcal{A}$ had not stopped by time $L$, we then define $M^{j}_{t,L}=\infty$ to ensure a lower bound on the original test statistic (if we could compute it). Therefore, for $L<\infty$,
@@ 2.2 Subtleties regarding Subroutine $\mathcal{L}$
For the proof of the upcoming theorem, we still need a definition of $M^{j}_{t,\infty}$ in the case that $L=\infty$ and $\tau_{j}=\infty$. We define
$M^{j}_{t,\infty}:=\begin{cases}M_{t}^{(\tau_{j}^{\prime})}(X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}),&\text{if }t\leq\tau_{j}^{\prime}<\infty,\\ -\infty,&\text{if }t>\tau_{j}^{\prime}\text{ or }\tau_{j}^{\prime}=\infty.\\ \end{cases}$ (2.4)
The rationale behind setting $M^{j}_{t,\infty}=-\infty$ when $\tau^{\prime}_{j}=\infty$ is the following. If $\tau_{j}=\infty$, for all $j=1,\cdots,B$, then we want to reject $H_{0,t}$, which is achieved by setting $M^{j}_{t,\infty}=-\infty$.
@@
The value of $L$ does not need to be a fixed constant; it can be set adaptively depending on the observed data sequence, for instance, one can choose $L$ to be $2\tau$. The proof of coverage works by arguing validity when $L=\infty$, and then pointing out that any smaller choice of $L$ only leads to more conservatism.
2.3 Prototypical choice of test statistic $M_{t}$
In the offline setting, where data is observed up to some fixed time $n$, the likelihood of data is
$L_{T}^{n}=\prod_{i=1}^{T-1}f_{0}(X_{i})\prod_{i=T}^{n}f_{1}(X_{i}),$ (2.5)
and the maximum likelihood estimator of $T$ [8, 26] is given by
$\hat{T}_{n}=\operatorname*{arg\,max}_{1\leq j\leq n}L_{j}^{n}=\operatorname*{arg\,max}_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$ (2.6)
In our sequential setting, where the data is observed up to a data-dependent stopping time $\tau$, we adapt this estimator by replacing $n$ with $\tau$ and define the estimator of the changepoint as:
The prototypical test statistic that we employ in our simulations is the likelihood-ratio-based test statistic $M_{t}^{(n)}$, where for all $n\in\mathbb{N},t\leq n$,
$M_{t}^{(n)}(X_{1},\cdots,X_{n}):=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}^{n}}{L_{t}^{n}}.$ (2.8)
Note that for all $t\in\mathbb{N}$, the test statistic $M_{t}$ for testing $H_{0,t}$ simplifies to
$M_{t}=\begin{cases}L_{\hat{T}}^{\tau}/L_{t}^{\tau},&\text{ if }t\leq\tau<\infty\\ -\infty&\text{if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (2.9)
2.4 Unconditional coverage when $\mathcal{A}$ controls PFA
We now investigate the coverage properties and provide a bound on the coverage of the confidence set produced by Algorithm 1.
@@ 2.4 Unconditional coverage when $\mathcal{A}$ controls PFA
Let $\mathcal{C}$ be the confidence set for changepoint $T\in\mathbb{N}$ produced by Algorithm 1, for some $\alpha\in(0,1)$, $B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$ and change detection algorithm $\mathcal{A}$. Then,
$\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha-\mathbb{\mathbb{P}}_{\infty}(\tau<T).$
Therefore, if $\mathcal{A}$ controls PFA at level $\delta$,
$\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha-\delta.$

### v3

@@ 4.1 Known pre- and post-change distributions
We develop a simulation-based scheme to approximate the exact $\alpha$-th quantile of the distribution of the test statistic $M_{t}$. Here, $M_{t}$ need not be of the form (3.4), it can be any test statistic based on $X_{1},\cdots,X_{\tau}$. Assuming known pre- and post-change distributions, we draw $B$ many independent streams of data under $H_{0,t}$. For the $j$-th sequence, denoted as $\{X_{n}^{j}\}_{n}$, we simulate data points until a change is detected at $\tau_{j}^{\prime}=\mathcal{A}(\{X_{n}^{j}\}_{n})$ or until a fixed time $L$, whichever happens first. Then, we compute the test statistic $M_{t,L}^{j}$ as described in Section 4.1.1, based on $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$. Our confidence set is the collection of all $t\in\{1,\cdots,\tau\}$, for which we fail to reject $H_{0,t}$:
@@
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-\alpha r_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})\right\},$ (4.1)
where $r_{t}$ is as defined in (3.7). Algorithm 1 contains an overview of the method.
@@
[…]
2:A confidence set $\mathcal{C}$ for the changepoint $T$
[…]
4:for $i=1$ to $N$ do
5:  Draw $\{Z_{n}^{i}\}_{n}\stackrel{{\scriptstyle}}{{\sim}}F_{0}$, until $\mathcal{A}$ detects at $\tau_{i}$, or until $\tau$, whichever happens first
6:end for
7:for $t=1$ to $\tau$ do
8:  Compute test statistic $M_{t}$ (e.g. (3.2)) based on $\tau$ and $X_{1},\dots,X_{\tau}$
9:  for $j=1$ to $B$ do
10:   Simulate $X_{1}^{j},\dots,X_{t-1}^{j}\stackrel{{\scriptstyle}}{{\sim}}F_{0}$; $X_{t}^{j},X_{t+1}^{j},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{1}$, until $\mathcal{A}$ detects a change at $\tau_{j}^{\prime}$, or time $L$, whichever happens first
11:   Subroutine $\mathcal{L}$: Compute $M_{t,L}^{j}$ based on $\tau_{j}^{\prime}$ and $X_{1}^{j},\dots,X_{\min(\tau_{j}^{\prime},L)}^{j}$
12:  end for
13:  if $M_{t}\leq\operatorname{Quantile}\left(1-\frac{\alpha}{N}\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t);\,M_{t},M_{t,L}^{1},\dots,M_{t,L}^{B}\right)$ then
14:   $\mathcal{C}\leftarrow\mathcal{C}\cup\{t\}$
15:  end if
16:end for
Theorem 4.1.
For any fixed $\alpha\in(0,1)$, $B,N,T\in\mathbb{N}$, ${}L\in\mathbb{N}\cup\{\infty\}$, and change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}_{1}$ defined in (4.1) satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-{\alpha}.$
Remark 4.2.
Notably, this method allows arbitrary (known) dependence among observations and is valid for any choice of $B,N,L$, test statistic $M_{t}$, and detection algorithm $\mathcal{A}$, which can function as an entirely black-box or heuristic method. The only necessity is a stopping rule that can be applied to generated data sequences.
4.1.1 Subtleties regarding Subroutine $\mathcal{L}$
The problem with $L=\infty$ is that the stopping time may never be reached for some sampled sequence. One can set $L=\infty$ if $\mathcal{A}$ is known to satisfy $\mathbb{P}_{T}(\tau<\infty)=1$, for all $T\in\mathbb{N}$ (if there is a change, $\mathcal{A}$ stops almost surely). This is a very mild assumption because any reasonable change detector should eventually raise an alarm if there is a change; indeed, for example, this holds for all algorithms with finite ARL. If $\mathcal{A}$ is not known to satisfy this condition, one can simply pick $L<\infty$. If $L<\infty$ and the $j$th sequence stopped because $\mathcal{A}$ raised an alarm at or before $L$, then we define $M^{j}_{t,L}$ as usual. If $\mathcal{A}$ had not stopped by time $L$, we then define $M^{j}_{t,L}=\infty$ to ensure a lower bound on the original test statistic (if we could compute it). For $L<\infty$,
@@ 4.1.1 Subtleties regarding Subroutine $\mathcal{L}$
We still need a definition of $M^{j}_{t,\infty}$ in the case that $L=\infty$ and $\tau_{j}=\infty$. We define $M^{j}_{t,\infty}:=M_{t}^{(\tau_{j}^{\prime})}(X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}),\text{if }t\leq\tau_{j}^{\prime}<\infty,$ and $M^{j}_{t,\infty}:=-\infty,\text{if }t>\tau_{j}^{\prime}\text{ or }\tau_{j}^{\prime}=\infty.$
@@
The value of $L$ does not need to be a fixed constant; it can be set adaptively depending on the observed data sequence, for instance, one can choose $L$ to be $2\tau$.
4.2 Adaptive confidence set for known pre-change distribution and parametric post-change class
We now assume that the post-change distribution follows some parametric composite model $\mathcal{P}_{1}=\{F_{\theta}:{\theta}\in\Theta_{1}\}$ and the pre-change distribution, $F_{\theta_{0}}(\theta_{0}\notin\Theta_{1})$ is known.
[…]
$\mathbb{P}(\forall n\in\mathbb{N},\theta\in\operatorname{CS}(X_{1},\cdots,X_{n};1-c))\geq 1-c.$ (4.3)
We construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, which have $1-\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$ coverage, where $r_{t}$ is as defined in (3.5). Despite $r_{t}$ being random, it turns out that marginalizing over $r_{t}$ yields that the CS has coverage $\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$; we formalize this more explicitly in the proof of the Theorem 4.4.
Fix $B\in\mathbb{N}$. For each $j=1,\cdots,B$, suppose we have a sequence $\{X^{j}_{n}(\theta_{0},t,\theta)\}_{n}$ $\sim\mathbb{P}_{\theta_{0},t,\theta}$. Note that for a fixed $t$ and $\theta$, the sequences must be independent across $j$. However, for any fixed $j$, the i.i.d. sequences $\{X^{j}_{n}(\theta_{0},\theta)\}_{n}$ are allowed to be dependent across different values of $t$ and $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$. A specific method for generating these sequences is detailed in Section B of the Supplementary. We draw the data until a change is detected at $\tau_{j,t}^{\theta}$ using $\mathcal{A}$ or until a fixed time $L$, whichever occurs first. Then, we compute the same test statistics truncated at $L$, based on $\tau_{j,t}^{\theta}$ and $X^{j}_{1}(\theta_{0},t,\theta),\cdots,X^{j}_{\tau_{j,t}^{\theta}\wedge L}(\theta_{0},t,\theta)$, which is denoted as $M^{j}_{t,L}(\theta)$. For $L<\infty$,
[…]
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\}$ (4.4)
Algorithm 2 summarizes the method. Note that, even when $\mathcal{S}^{\prime}_{\tau-t+1}$ is an infinite set, the threshold $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ can be approximated by discretizing $\mathcal{S}^{\prime}_{\tau-t+1}$ into a grid and evaluating the quantile for values of $\theta$ on this grid, taking their maximum. While this approach is conceptually straightforward, it is computationally intensive; related experiments are in Section E of the Supplementary. Moreover, in many cases, we can derive a suitable upper bound of this threshold, which is computationally, statistically efficient and does not require any approximation. Details are provided in Section B of the Supplementary, along with two concrete examples of the Gaussian and Poisson mean change problems.
Algorithm 2 CI for known pre-change ($F_{\theta_{0}}$) and parametric post-change
[…]
2:Confidence sets $\mathcal{C}$ for changepoint $T$
3:for $i=1,\cdots,N$ do
4:  Draw $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta_{0}}$, until $\mathcal{A}$ detects at $\tau_{i}$ or until $\tau$, whichever happens first.
5:end for
[…]
7:for $t=1,\cdots,\tau$ do
8:  Compute some test statistics $M_{t}$ based on $X_{1},\cdots,X_{\tau}$
9:  Compute $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$
10:  Find a confidence sequence for $\theta_{1}$: $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$
11:  for $j=1,\cdots,B$ do
12:   Subroutine 2:
13:   for $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}(X_{t},\cdots,X_{\tau})$ do
14:     Get $\{X^{j}_{n}(\theta_{0},t,\theta)\}_{n}$ following $\mathbb{P}_{\theta_{0},t,\theta}$ (independently for all $j$), until its detection time $\tau_{j,t}^{\theta}$ using $\mathcal{A}$ or until time $L$, whichever happens first.
15:     Compute $M_{t,L}^{j}(\theta)$ based on $\tau_{j,t}^{\theta}$, and $X^{j}_{1}(\theta_{0},t,\theta),\cdots,X^{j}_{\tau_{j,t}^{\theta}\wedge L}(\theta_{0},t,\theta)$
16:   end for
17:  end for
18:  Subroutine 3:
19:  if there exists $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ then
20:   $\mathcal{C}=\mathcal{C}\cup\{t\}$
21:  end if
22:end for
@@ 4.2 Adaptive confidence set for known pre-change distribution and parametric post-change class
Let $\mathcal{C}_{1}$ be the confidence set for changepoint $T\in\mathbb{N}$ as defined in (4.4) for some $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N}$, ${}L\in\mathbb{N}\cup\{\infty\}$. Then for any given detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$ and any $\theta_{1}\in\Theta_{1}$, $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta.$
4.3 Adaptive confidence set for parametric pre- and post-change classes
We now assume that both pre- and post-change distributions follow some parametric composite models: $\mathcal{P}_{0}=\{F_{\theta}:\theta\in\Theta_{0}\}$ and $\mathcal{P}_{1}=\{F_{\theta}:\theta\in\Theta_{1}\}$, $\Theta_{0}\cap\Theta_{1}=\emptyset$.
[…]
For each $j=1,\cdots,B$, for some fixed $B\in\mathbb{N}$, suppose we have sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ having joint distribution as $\mathbb{P}_{\theta,t,\theta^{\prime}}$. Note that the sequences must be independent as $j$ varies. However, for any fixed $j$, the i.i.d. sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ can be dependent across $t$ and $\theta\in\mathcal{S}_{t-1},\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$. A specific method for generating these sequences is detailed in Section C of the Supplementary. We draw the data until a change is detected at the stopping time $\tau^{j,t}_{\theta,\theta^{\prime}}$ using $\mathcal{A}$ or until time $L$, whichever occurs first. Then, we compute the same test statistics truncated at $L$, based on $\tau^{j,t}_{\theta,\theta^{\prime}}$ and $X^{j}_{1}(\theta,t,\theta^{\prime}),\cdots,X^{j}_{\tau^{j,t}_{\theta,\theta^{\prime}}\wedge L}(\theta,t,\theta^{\prime})$, which is denoted as $M^{j}_{t,L}(\theta,\theta^{\prime})$. Similar to (4.4), the adaptive confidence set is defined as
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{\mathbf{S}}^{(t)}}\operatorname{Quantile}(1-{\alpha}r_{t}^{*};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}$ (4.5)
where $\mathcal{\mathbf{S}}^{(t)}=\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. Algorithm 3 contains an overview. If the detection algorithm controls PFA, a simpler algorithm can be used, which avoids Assumption 3.3 and provides coverage guarantees, as elaborated in Section F of the Supplementary.
As before, one can always adopt a grid-based approach, which is conceptually straightforward, but it is computationally intensive; corresponding experimental results are reported in Section E of the Supplementary. Additionally, in many cases, we can derive a suitable upper bound of this threshold, which is computationally efficient and exact, but that can be a bit more conservative. We provide the details in Section C of the Supplementary, along with two concrete examples.
Algorithm 3 CI for parametric pre-change and post-change
[…]
2:A confidence set $\mathcal{C}$ for changepoint $T$
[…]
4:for $i=1$ to $N$ do
5:  Draw $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F_{\theta_{0}^{*}}$, until $\mathcal{A}$ detects at $\tau_{i}^{*}$ or until $\tau$, whichever happens first.
6:end for
7:for $t=1$ to $\tau$ do
8:  Compute some test statistics $M_{t}$ based on $X_{1},\cdots,X_{\tau}$
9:  Compute $r_{t}^{*}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}^{*}\geq t)/N$
10:  Find a confidence interval for $\theta_{0}$: $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t}^{*})$
11:  Find a confidence sequence for $\theta_{1}$: $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t}^{*})$
12:  for $j=1$ to $B$ do
13:   Subroutine 1:
14:   for each $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ do
15:     Simulate $\{X_{n}^{j}(\theta,t,\theta^{\prime})\}_{n}$ following $\mathbb{P}_{\theta,t,\theta^{\prime}}$ (independently for all $j$) until its detection time $\tau^{j,t}_{\theta,\theta^{\prime}}$ using $\mathcal{A}$ or until time $L$, whichever happens first
16:     Compute $M_{t,L}^{j}(\theta,\theta^{\prime})$ based on $\tau^{j,t}_{\theta,\theta^{\prime}}$, $X_{1}^{j}(\theta,t,\theta^{\prime}),\cdots,X_{\tau^{j,t}_{\theta,\theta^{\prime}}\wedge L}^{j}(\theta,t,\theta^{\prime})$
17:   end for
18:  end for
19:  Subroutine 2:
20:  if $M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\alpha r_{t}^{*};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$ then
21:   $\mathcal{C}\leftarrow\mathcal{C}\cup\{t\}$
22:  end if
23:end for
Theorem 4.5.
Let $\mathcal{C}_{1}$ be the confidence sets for changepoint $T\in\mathbb{N}$ defined in (4.5), for some $\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N}$, ${}L\in\mathbb{N}\cup\{\infty\}$. For any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1}$, and any change detection algorithm $\mathcal{A}$ satisfying Assumption 3.3 and having $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, the following guarantee holds: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma.$